OpenStudy (anonymous):

What is the rate of change for f(x) = 4 sin x − 2 on the interval from x = 0 to x = pi over 2?

OpenStudy (anonymous):

pi over 4 4 over pi pi over 8 8 over pi

OpenStudy (anonymous):

@Peaches15

jimthompson5910 (jim_thompson5910):

if you replace x with 0, what is the value of f(x) ?

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 4\sin(x) - 2\] \[\Large f(0) = 4\sin(0) - 2\] \[\Large f(0) = ???\]